OpenStudy (anonymous):

. A car drives 12 miles north, turns, and then drives 8 miles east. How far is the car from its starting point? a. Draw a right triangle to represent this situation. Label the lengths of the known sides. Label the length of the unknown side x.

OpenStudy (whpalmer4):

Draw your picture... put an x on the right spot Use Pythagorean theorem to find the value of \(x\): \[a^2+b^2=x^2\]\[12^2+8^2=x^2\]\[x=\]

OpenStudy (anonymous):

thanks i needed to check my answer